# Public Projects with Preferences and Predictions

### v1

@@ A.2 Two Alternatives
In the rest of this section, we prove Theorem 1,. We first observe an equivalent representation of the first-order conditions in the two-alternative case.
@@
[…]
@@
Proof of Proposition 2.
@@
On the other hand, we can replace Inequality 33 as follows:
@@ A.3 QTM with m alternatives
Here we prove Proposition 3, that $\text{pPoA}\geq\frac{1}{m}$ for the QTM.
Proof of Proposition 3.
@@
Appendix B Wagering Mechanisms
Lemma (Lemma 6 restated).
The importance-weighted quadratic decision wagering mechanism with parameter $\beta=\epsilon x$ has a deviation bound $\epsilon^{1/2}x$.
@@ Appendix B Wagering Mechanisms
Consider any participant $i$. The only terms in their expected utility that $i$ controls are $\frac{1}{p_{k}}s(\hat{b}_{k}^{i},b_{k}^{*})$ for each $k$. Exactly as in the prediction-market analysis , the expected utility of this term is given by Equation 10, $S(\mathbf{\hat{b}}^{i},\mathbf{b}^{i})$, where $\mathbf{b}^{i}$ is their belief. (The same comments about the variance term apply.) By the same arguments as in Lemma 5, with parameter $\beta=\epsilon x$, the loss from misreporting is
$S(\mathbf{\hat{b}}^{i},\mathbf{b}^{i})-S(\mathbf{b}^{i},\mathbf{b}^{i})=\frac{-1}{\epsilon x}\sum_{k}(\hat{b}^{i}_{k}-b^{i}_{k})^{2}.$
If the loss is bounded by $x$, then $|\hat{b}^{i}_{k}-b^{i}_{k}|\leq\epsilon^{1/2}x$.
Now, this holds for all participants, and the output of the wagering mechanism is $\hat{B}_{k}=\frac{1}{N}\sum_{i=1}^{N}\hat{b}^{i}_{k}$, so the mechanism has a deviation bound $\epsilon^{1/2}x$. ∎

### v2

@@ A.2 Two Alternatives
We begin by proving Theorem 1,. We first observe an equivalent representation of the first-order conditions in the two-alternative case.
@@
In any pure-strategy equilibrium of the two-alternative QTM, $p_{1}\geq 1/2$, with strict inequality if $V_{1}>V_{2}$. In particular, if $c\geq\frac{\max_{i,k}v_{k}^{i}}{2}$, then $\text{pPoA}(\mathbf{v})>1/2$.
@@
Total payments.
Proposition 7.
In the QTM with two alternatives and type profile $\mathbf{v}$, with $V_{1}>V_{2}$, define the disagreement to be $D:=\frac{\sum_{i}(v_{1}^{i}-v_{2}^{i})^{2}}{(V_{1}-V_{2})^{2}}$. Then in any pure-strategy equilibrium, [^4] Here $\Theta(\cdot)$ is with respect to $V_{1}-V_{2}$ growing, or at least bounded below. The proof includes fully explicit nonasymptotic bounds.
$\text{revenue}=\Theta\left(c~{}D\left(\ln\tfrac{V_{1}-V_{2}}{c}\right)^{2}\right).$
Proof.
@@
On the other hand, we can replace Inequality 19 as follows:
@@
  • For $n_{1}-n_{2}\ll\sqrt{n}$, $D=\omega(1)$ and revenue grows more rapidly.
@@ A.3 QTM with m alternatives
We briefly touch on the case where there are more than two alternatives. Eguia et al. (2019, 2023) give asymptotic results showing that in their settings, with a fixed distribution of bounded agents, as $n\to\infty$ the welfare of the QTM tends to optimality. Here we can give a weaker but nonasymptotic bound of $\geq\frac{1}{m}$. The proof is not immediate and involves dividing the alternatives into a group with “high” welfare (higher than the mechanism’s expectation) and the remainder, an idea from Eguia et al. (2023) used in a different way.
Proof of 2.
@@
It is not clear if this bound can be improved non-asymptotically. The analysis of the many-alternatives case is very complex (see Eguia et al. (2019, 2023)). As far as we currently know, it may be possible to have instances of a large “spread” $\frac{V_{1}}{\max_{i,k}v_{k}^{i}}$, perhaps exponentially large in $m$, with pure-strategy Price of Anarchy bounded away from one, or perhaps as small as $O(1/m)$.
Appendix B SQUAP Price of Anarchy
Proof of Theorem 3.
By Proposition 3, we have $\min_{\mathbf{p}\in\text{pPNE}(\mathbf{v},\mathcal{D})}\frac{\sum_{k}p_{k}\hat{W}_{k}}{\hat{W}_{1}}\geq C(\mathbf{v})$, and we know $\max_{k}|\hat{B_{k}}-B_{k}|\leq\alpha\max_{i,k}v_{k}^{i}$ by Lemma 1. It follows that, for $\mathbf{p}\in\arg\min_{\mathbf{p}\in\text{pPNE}(\mathbf{v},\mathcal{D})}\frac{\sum_{k}p_{k}W_{k}}{W_{1}}$,
$\text{pPoA}(\mathbf{v},\mathcal{D})$ $=\frac{\sum_{k}p_{k}W_{k}}{W_{1}}$
$\geq\frac{\sum_{k}p_{k}\hat{W}_{k}-\alpha\max_{i,k}v_{k}^{i}}{\hat{W}_{1}+\alpha\max_{i,k}v_{k}^{i}}$
$\geq\frac{\sum_{k}p_{k}\hat{W}_{k}-\alpha\max_{i,k}v_{k}^{i}}{\hat{B}_{1}+(1+\alpha)V_{1}}$
$\geq\frac{\sum_{k}p_{k}\hat{W}_{k}-\alpha\max_{i,k}v_{k}^{i}}{(1+\alpha)\hat{W}_{1}}$
$\geq\frac{1}{1+\alpha}\left(C(\mathbf{v})-\frac{\alpha}{T}\right),$
where the second inequality follows from the fact that $\alpha\max_{i,k}v_{k}^{i}\leq\alpha V_{1}$. ∎
Appendix C Prediction Markets
We prove the $(\alpha,x)$-robustness of the importance-weighted decision market.
@@ Appendix C Prediction Markets
Observe that under the efficient market assumption, the decision market is equivalent to the following: run the entire market, where the last participant predicts $\mathbf{B^{*}}$. Then, give the last participant one more opportunity to participate; denote their final predictions by $\mathbf{\hat{B}}$. This is equivalent because of the telescoping sum of the market scoring rule, so that the two prediction opportunities have the same net payoff as one opportunity with predicting $\mathbf{\hat{B}}$.
For any participant except the final one, under the efficient market assumption, their strategy does not affect the final predictions $\mathbf{\hat{B}}$. One way to see this is by the above equivalence, since the market will end at $\mathbf{B}^{*}$ regardless, followed by the last participant’s possible manipulation.
So we only need to analyze the incentives of the final participant $i$. By the decomposition above, we only need to consider the incentives for the final prediction opportunity (since the previous one is fixed to $\mathbf{B^{*}}$). The expected utility for $\mathbf{\hat{B}}$ is
$S(\mathbf{\hat{B}};\mathbf{B^{*}})-S(\mathbf{B^{*}};\mathbf{B^{*}})$ $=S(\mathbf{\hat{B}};\mathbf{B^{*}})$
$=-\frac{1}{\beta}\sum_{k}(\hat{B}_{k}-B_{k}^{*})^{2}.$ (via Equation 7)
By the definition of $x$-best-responding in the market, $x>\frac{1}{\beta}\sum_{k}(\hat{B}_{k}-B_{k}^{*})^{2}$. So for all $k$, $(\hat{B}_{k}-B_{k}^{*})^{2}\leq\beta x=\epsilon x^{2}$, or $|\hat{B}_{k}-B_{k}^{*}|\leq\sqrt{\epsilon}x$. ∎
We prove that, in strategy profiles with high “disagreement”, SQUAP with prediction markets is budget balanced in expectation.
Proof of Corollary 5.
We prove that the mechanism can choose liquidity parameters $\beta$ so that the expected revenue is larger than the amount spent on the decision market stage in expectation. In practice, one would like to scale the prediction-market payments by the largest possible amount so that budget balance is achieved. However, we cannot rule out manipulation for reasons similar to the redistribution, so for analysis we assume that the mechanism designer commits to $\beta$ up front. By the telescoping nature of the proper scoring rule, the mechanism’s total spend on the decision market stage in expectation is $S(\mathbf{\hat{B}};\mathbf{B^{*}})-S(\mathbf{B^{0}},\mathbf{B^{*}})\leq S(\mathbf{B^{*}};\mathbf{B^{*}})-S(\mathbf{B^{0}};\mathbf{B^{*}})=\frac{1}{\beta}\sum_{k}(B_{k}^{0}-B_{k}^{*})^{2}$. With at-least-constant disagreement, revenue is at least constant (by Proposition 7), so $\beta=\Omega(1)$ is possible with ex ante budget-balance. By Corollary 4, in this case and with $T\to\infty$, the Price of Anarchy converges to $1$. ∎